The demand curve for the output of a certain industry is linear; q = A - Bp. There are constant marginal costs of C. For all values of A, B, and C such that A > 0, B > 0, and 0 < C < A/B,

Group of answer choices

none of the other answers are correct

if the industry is monopolized, prices will be exactly twice as high as they would be if the industry were competitive.

if the industry is monopolized, output will be more than half as large as it would be if the industry were competitive.

if the industry is competitive, output will be exactly twice as great as it would be if the industry were monopolized.

if the industry is monopolized, prices will be more than twice as high as if the industry is competitive.
Answer #1

Option (3).

q = A - Bp

Bp = A - q

p = (A - q) / B

In perfect competition, p = MC.

(A - q) / B = C

A - q = BC

q = A - BC

p = [A - (A - BC)] / B = BC / B = C

In monopoly, MR = MC.

TR = p x q = (Aq - q2) / B

MR = dTR/dq = (A - 2q) / B

(A - 2q) / B = C

A - 2q = BC

2q = A - BC

q = (A - BC) / 2 [Monopoly output = (1/2) x Competitive output, i.e. Competitive output = 2 x Monopoly output]